OpenStudy (anonymous):

What is the solution to the equation 4(2b - 4) = 2(3 - 5b)

OpenStudy (hba):

See this a(b+c) =)a)(b)+(a)(c)

OpenStudy (anonymous):

When you do an equation like this, you should distribute the 4 outside of the brackets onto the (2b-4). You should do the same with the other side of the equation. If you do this, do you know what your equation should look like now or should I explain further?

OpenStudy (anonymous):

A. b = \[\frac{ 7 }{ 11 }\] B. b = \[\frac{ 9 }{ 11 }\] C. b = \[\frac{ 11 }{ 9 }\] D. b = \[\frac{ 11 }{ 7 }\]

OpenStudy (hba):

Try solving it on your own.

OpenStudy (anonymous):

yea thats what i am doing

OpenStudy (anonymous):

b= \[\frac{ 11 }{ 9 }\] so the answer is c

OpenStudy (hba):

Yeah correct :)

OpenStudy (anonymous):

tx